OpenStudy (anonymous):

Can x-3/x+3 be simplified further?

OpenStudy (zpupster):

that should be it

Nnesha (nnesha):

if it's \[\huge\rm \frac{ x-3 }{ x}+3\] then you can simplify

Nnesha (nnesha):

if yes:then 3 is same as 3/1 \[\huge\rm \frac{ x-3 }{ \color{blue}{x} }+\frac{ 3 }{ \color{ReD}{1}}\] the common denominator is x and when we find the common denominator we should multiply the numerator of `1st` fraction with the denominator of `2nd` fraction and multiply the numerator of `2nd` fraction by the denominator of `1st` fraction \[ \huge\rm \frac{\color{reD}{1}( x-3) +3 \times \color{blue}{x}}{ x } \] now simplify the numerator feel free to tag me if u have any question

OpenStudy (anonymous):

No it's just (x-3)/(x+3)

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

then what zpupster said.

Nnesha (nnesha):

do you have options (answer choices )?

Nnesha (nnesha):

thats already simplified. you can divide them using long or synthetic division

OpenStudy (anonymous):

@Nnesha No options. Okay. Thank you!

Nnesha (nnesha):

alright then